OpenStudy (anonymous):

Multiply or divide. Show your work. 5x+10/x+2 * 2x/4x-10

OpenStudy (anonymous):

\[\frac{ a }{ b } \times \frac{ c }{ d } \implies \frac{ a \times c }{ b \times d }\]

OpenStudy (amorfide):

do as @iambatman said, then you can simplify

OpenStudy (anonymous):

5x/20 * 8x/-10

OpenStudy (anonymous):

i'm not sure if is that right

Nnesha (nnesha):

\[\frac{ 5x+10 }{ x+2 } \times \frac{ 2x }{ 4x-10 }\] first factor 5x+10 and 4x-10

OpenStudy (anonymous):

what I should do Multiply or divide.

OpenStudy (anonymous):

5x^2+20

Nnesha (nnesha):

okay for multiplication you have to multiply everything by 2x so 2x times 5x = ??

OpenStudy (anonymous):

10x

Nnesha (nnesha):

when you multiply variables you should add their exponent

OpenStudy (anonymous):

oh ok 10x^2

Nnesha (nnesha):

yep right now multiply denominator by denominator

OpenStudy (anonymous):

im not sure wat to Multiply next

Nnesha (nnesha):

okay  first multiply numerator